OpenStudy (anonymous):

the greatest acceleration or deceleration that a train may have is a.The minimum time in which the train can start from one station to stop at the next is a. 2sqrt{s/a} b. sqrt{s/a}

OpenStudy (anonymous):

As in the figure, for the train to take minimum time, it should accelerate with it's max acceleration for half the duration and decelerate for the next half. $$s = \frac12 t v_{max} = \frac12 t(\frac{at}{2}) $$ It gives $$ t= 2\sqrt{\frac sa}$$

OpenStudy (anonymous):

tnx a lot.. :)

OpenStudy (anonymous):

also try the v-t (velocity - time) graph method...in complex cases it is really useful...and more easy to understand.just have to keep 2 things in mind: 1)equate the slope of graph to acceleration. 2)equate the area under graph to total displacement....

OpenStudy (anonymous):

thank u.. :)